OpenStudy (anonymous):

\[6x \over 5y 10y \over8x\]

OpenStudy (angela210793):

6x/5y 10y/8x...wht do u do with these two?

OpenStudy (anonymous):

I don't know, it is missing an operator(+, -, *, /, etc.)

OpenStudy (anonymous):

multipy together

OpenStudy (angela210793):

=60xy/40xy=6/4=3/2